Solve the problem. Round the point estimate to the nearest thousandth.

32 randomly picked people were asked if they rented or owned their own home, 8 said they rented. Obtain a point estimate of the proportion of home owners.
Answer #1

Solution :

Given that,

n = 32

x = 8

Point estimate = sample proportion = = x / n = 8/32=0.25